OpenStudy (anonymous):

Which value is the least? the mixed number one and one-fourth the mixed number one and one-third the decimal one and two tenths the decimal one and three tenths Which value is the least? the mixed number one and one-fourth the mixed number one and one-third the decimal one and two tenths the decimal one and three tenths

jimthompson5910 (jim_thompson5910):

convert every number to a decimal and that will allow you to sort/compare them

jimthompson5910 (jim_thompson5910):

for example 1 and 1/4 = 1 + 1/4 1 and 1/4 = 1 + 0.25 1 and 1/4 = 1.25

jimthompson5910 (jim_thompson5910):

1 and 1/3 = 1 + 1/3 1 and 1/3 = 1 + 0.333333 1 and 1/3 = 1.333333 Do this for all of the mixed numbers and fractions given

OpenStudy (anonymous):

ohhh i get it know thank you very much :)

jimthompson5910 (jim_thompson5910):

yw